OpenStudy (anonymous):

I really need help coming up with the equation to this problem. Help please ! A snail oozes at a uniforn1 speed along the crossbar of a picket fence. On day zero, it is at the third picket; on day five, it has reached the tenth picket. Write a linear model of the snail's position on day n. In what units is the slope measured?

OpenStudy (anonymous):

units of slope : fenceposts/day

OpenStudy (anonymous):

day zero, third picket (x, y) = (0, 3) day five, tenth picket (x, y) = (5, 10) slope = (y^2-y^1) / (x^2-x^1) (10-3) / (5-0) slope = 7/5 To solve for the equation of the line, you can use either y = mx + b (slope-intercept form) or (y-y1) = m (x-x1) (point-slope form).

OpenStudy (anonymous):

Thanks so much !